OpenStudy (anonymous):

Use the definition of the derivative to find (2/[sqrt(x)] )'.

OpenStudy (isaiah.feynman):

\[f'(x)=\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h } \]

OpenStudy (isaiah.feynman):

\[f'(x) =\lim_{h \rightarrow0} \frac{ \frac{ 2 }{ \sqrt{x+h} }-\frac{ 2 }{ \sqrt{x} } }{ h }\]